# Insights into the Lottery Ticket Hypothesis and Iterative Magnitude Pruning

### v2

@@ V Conclusion and Scope for Further Work
TABLE XI: Comparison of $\lambda_{max}$ and $V^{{}^{\prime}}(100)$ at $W^{RPr{(min\_(L))}}_{(L-1)}$ and $W_{(L-1)}^{(min\_(L-1))}$ for $L=$$1,6,12$} in case of VGG-16.
$L$ Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$
1 $W^{RPr{(min\_(1))}}_{(0)}$ 43.777 219.996
$W_{(0)}^{(min\_(0))}$ 67.929 220.915
6 $W^{RPr{(min\_(6))}}_{(5)}$ 46.733 201.835
$W_{(5)}^{(min\_(5))}$ 77.396 189.044
12 $W^{RPr{(min\_(12))}}_{(11)}$ 4379.201 426.294
$W_{(11)}^{(min\_(11))}$ 1326.092 352.86
@@
Fig. 28 presents the training loss along a straight line connecting $W^{(min\_(11))}_{(11)}$ (baseline for $W^{(RIPN)}$) and $W^{(RIPN)}$. The figure shows a huge barrier between the two points, which demonstrates that these two points lie in different loss sublevel sets.
Figure 28: Training Loss along a straight line between $W^{(min\_(11))}_{(11)}$ and $W^{(RIPN)}$.
A comparison of top-100 positive eigen values of the Hessian at $W^{(RIPN)}$ and $W^{(min\_(12))}_{(12)}$ is given in Fig. 29. It can be observed from the figure that the Hessian at $W^{(RIPN)}$ has larger eigen values than that at $W^{(min\_(12))}_{(12)}$, which indicates a smaller volume for the basin around $W^{(RIPN)}$ (Table XII).
Figure 29: Comparison of top-100 positive eigen values of the Hessian at $W^{(RIPN)}$ and $W^{(min\_(12))}_{(12)}$.
TABLE XII: Comparison of $\lambda_{max}$ and $V^{{}^{\prime}}(100)$ at $W^{(RIPN)}$ and $W^{(min\_(12))}_{(12)}$
. Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$ $W^{(RIPN)}$ 25526.326 442.242 $W^{(min\_(12))}_{(12)}$ 3992.4783 389.957
@@
Fig. 30 presents the training loss along a straight line connecting $W^{(min\_(11))}_{(11)}$ (baseline for $W^{(RPN)}$) and $W^{(RPN)}$. The figure shows a huge barrier between the two points, which demonstrates that these two points lie in different loss sublevel sets.
Figure 30: Training Loss along a straight line between $W^{(min\_(11))}_{(11)}$ and $W^{(RPN)}$.
A comparison of top-100 positive eigen values of the Hessian at $W^{(RPN)}$ and $W^{(min\_(12))}_{(12)}$ is given in Fig. 31. It can be observed from the figure that the Hessian at $W^{(RPN)}$ has larger eigen values than that at $W^{(min\_(12))}_{(12)}$, which indicates a smaller volume for the basin around $W^{(RIPN)}$ (Table XIII).
Figure 31: Comparison of top-100 positive eigen values of the Hessian at $W^{(RPN)}$ and $W^{(min\_(12))}_{(12)}$.
TABLE XIII: Comparison of $\lambda_{max}$ and $V^{{}^{\prime}}(100)$ at $W^{(RPN)}$ and $W^{(min\_(12))}_{(12)}$
Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$
$W^{(RPN)}$ 51190.094 451.384
$W^{(min\_(12))}_{(12)}$ 3992.4783 389.957

### v3

@@ V Conclusion and Scope for Further Work
TABLE VII: Comparison of $V^{{}^{\prime}}(100)$ at $W^{RPr{(min\_(L))}}_{(L-1)}$ and $W_{(L-1)}^{(min\_(L-1))}$ for $L=$$1,6,12$} in case of VGG-16.
$L$ Solution $V^{{}^{\prime}}(100)$
1 $W^{RPr{(min\_(1))}}_{(0)}$ 219.996
$W_{(0)}^{(min\_(0))}$ 220.915
6 $W^{RPr{(min\_(6))}}_{(5)}$ 201.835
$W_{(5)}^{(min\_(5))}$ 189.044
12 $W^{RPr{(min\_(12))}}_{(11)}$ 426.294
$W_{(11)}^{(min\_(11))}$ 352.86
@@
Fig. 29 presents the training loss along a straight line connecting $W^{(min\_(11))}_{(11)}$ (baseline for $W^{(RIPN)}_{(12)}$) and $W^{(RIPN)}_{(12)}$. The figure shows a huge barrier between the two points, which demonstrates that these two points lie in different loss sublevel sets.
Figure 29: Training Loss along a straight line between $W^{(min\_(11))}_{(11)}$ and $W^{(RIPN)}_{(12)}$.
A comparison of top-100 positive eigen values of the Hessian at $W^{(RIPN)}_{(12)}$ and $W^{(min\_(12))}_{(12)}$ is given in Fig. 30. It can be observed from the figure that the Hessian at $W^{(RIPN)}_{(12)}$ has larger eigen values than that at $W^{(min\_(12))}_{(12)}$, which indicates a smaller volume for the basin around $W^{(RIPN)}_{(12)}$ (Table VIII).
Figure 30: Comparison of top-100 positive eigen values of the Hessian at $W^{(RIPN)}_{(12)}$ and $W^{(min\_(12))}_{(12)}$.
TABLE VIII: Comparison of $V^{{}^{\prime}}(100)$ at $W^{(RIPN)}_{(12)}$ and $W^{(min\_(12))}_{(12)}$
. Solution $V^{{}^{\prime}}(100)$ $W^{(RIPN)}_{(12)}$ 442.242 $W^{(min\_(12))}_{(12)}$ 389.957
@@
Fig. 31 presents the training loss along a straight line connecting $W^{(min\_(11))}_{(11)}$ (baseline for $W^{(RPN)}_{(12)}$) and $W^{(RPN)}_{(12)}$. The figure shows a huge barrier between the two points, which demonstrates that these two points lie in different loss sublevel sets.
Figure 31: Training Loss along a straight line between $W^{(min\_(11))}_{(11)}$ and $W^{(RPN)}_{(12)}$.
A comparison of top-100 positive eigen values of the Hessian at $W^{(RPN)}_{(12)}$ and $W^{(min\_(12))}_{(12)}$ is given in Fig. 32. It can be observed from the figure that the Hessian at $W^{(RPN)}_{(12)}$ has larger eigen values than that at $W^{(min\_(12))}_{(12)}$, which indicates a smaller volume for the basin around $W^{(RPN)}_{(12)}$ (Table IX).
Figure 32: Comparison of top-100 positive eigen values of the Hessian at $W^{(RPN)}_{(12)}$ and $W^{(min\_(12))}_{(12)}$.
TABLE IX: Comparison of $V^{{}^{\prime}}(100)$ at $W^{(RPN)}_{(12)}$ and $W^{(min\_(12))}_{(12)}$
Solution $V^{{}^{\prime}}(100)$
$W^{(RPN)}_{(12)}$ 451.384
$W^{(min\_(12))}_{(12)}$ 389.957